# AAA-Gaussians: Anti-Aliased and Artifact-Free 3D Gaussian Rendering

### v1

@@ 1 Introduction
3D Gaussian Splatting (3DGS) [13] has recently revolutionized inverse rendering by enabling fast, differentiable rasterization of 3D Gaussian point clouds. While 2D splat evaluation is highly efficient, the projection from 3D to 2D Gaussians remains an approximation, introducing artifacts that become particularly noticeable under non-standard camera settings, such as a wide field-of-view in virtual reality rendering. Additionally, 3DGS further approximates the rendering of 2D splats, by assuming them to be parallel to the current view-plane, leading to blend order inconsistencies and popping artifacts under simple camera rotations [27].
@@
Several works attempt to bridge the gap between 2D splatting and 3D ray tracing by first bounding 3D Gaussians in screen space and then computing their contributions per ray in 3D [27, 36, 10, 32, 31]. While these hybrid 2D/3D approaches address certain limitations of 3DGS, their reliance on screen space computations still makes them prone to artifacts, particularly when rendering out-of-distribution camera poses—i.e., viewpoints or parameters significantly different from the training data (cf. Fig. 1): (1) Even when considering the highest contribution point in 3D [36, 27] or adjusting transformations per Gaussian [32], distortions may still occur. (2) Zooming in or out beyond the typical training views introduces artifacts, especially when evaluating Gaussians in 3D, as 2D anti-aliasing techniques can no longer be applied [35]. (3) Despite the use of 3D bounding planes, screen space computations become unstable when Gaussians extend behind the image plane, leading to artifacts at image boundaries [10]. (4) Simplified per-pixel sorting strategies that focus on high-opacity Gaussians, can result in inaccurate renderings, particularly for viewpoints far from the training distribution [10].
@@
Overall, our method achieves state-of-the-art quality in novel view synthesis, while allowing for artifact-free rendering in real-time.
@@ 2 Related Work
In this section, we cover recent radiance field and 3D Gaussian Splatting methods, with a focus on artifact-free rasterization and ray-tracing of 3D Gaussian representations.
@@ 2.1 Radiance Fields & 3D Gaussian Splatting
Radiance fields have attracted widespread interest in novel view synthesis since the publication of Neural Radiance Fields (NeRF) [21], which uses large, coordinate-based MLPs to query view-dependent color and density at any point in a bounded domain. Follow-up work includes improvements in terms of anti-aliasing [1, 3], extending NeRFs to unbounded scenes [2, 25], more efficient encodings [24], or fast rendering [9, 6, 30]. Despite the aforementioned improvements, most NeRF methods still require multiple costly MLP evaluations for each pixel, manifesting in long training and rendering times.
More recently, 3D Gaussian Splatting [13] exploded in popularity, replacing the slower implicit representations of NeRFs with an explicit 3D Gaussian point cloud representation that can be efficiently rasterized through elliptical weighted average (EWA) splatting [38]. The initial sparse point cloud can either be initialized randomly, from Structure-from-Motion [28], or from other guidances (e.g., sampled from a pretrained NeRF [26]). To achieve good coverage of the scene, this point cloud has to be densified by adding, cloning, or pruning points based on screen space gradients [13], per-view saliency maps [19], depth supervision [14], or by relocating low-opacity Gaussians [15]. Additionally, strategies such as opacity decay [27] and iterative pruning [8, 7] are employed to reduce the size of the resulting point cloud.
@@ 2.2 Artifacts in 3D Gaussian Splatting
The original 3D Gaussian Splatting suffers from a number of artifacts (which we underline), most of which were addressed in recent related work. Aliasing artifacts due to undersampling of the rasterized Gaussians can be solved by approximating the integral over each pixel [17], using multi-scale 3D Gaussians [34], or by applying a smoothing filter to both the 3D Gaussian and its projected 2D splat [35]. Popping artifacts occur due to the global sort of primitives before rasterization, leading to sudden changes in the blending order during view rotation. This classical problem of rasterizing semi-transparent surfaces can be solved by employing order-independent transparency methods [33], but requires accurate per-pixel sorting and depth values. StopThePop [27] computes Gaussian depth values along each view ray and uses hierarchical $k$-buffers to improve sort order. Other recent work relies on hybrid transparency [20], where only important contributors are sorted, while low-opacity Gaussians are blended in a "tail" [10]. Even perfect per-pixel sorting would lead to blending artifacts, as it assumes that primitives are non-overlapping and approximates them as surfaces, however, this cannot be solved analytically for Gaussians [18] and requires expensive volumetric integration techniques [5]. Furthermore, projection artifacts occur due to the affine approximation when projecting 3D Gaussians to 2D splats, resulting in disturbing cloud-like artifacts and elongated Gaussians at the image border. These artifacts are especially pronounced in large field of view settings, e.g. in virtual reality (VR). Recent methods solve this by projecting Gaussians onto the unit sphere tangent plane [12] or using Unscented Transform [32]. The easiest way to circumvent projection artifacts is to evaluate Gaussians in 3D. Several methods use ray tracing for this purpose [23, 18, 4], however, these methods are computationally expensive and require additional acceleration data structures, as well as specialized ray tracing hardware to achieve competitive performance. To circumvent the in-order ray tracing overhead, several methods propose to bound the Gaussian on screen, and then evaluate it in 3D by finding its point of maximum contribution along the ray. However, most methods rely on either the inaccurate 2D bounds from the approximate affine projection [36, 27, 31] or Unscented Transform [32]. Only recently, Hahlbohm et al. [10] proposed to perform plane fitting to find accurate screen space bounds of the 3D Gaussian ellipsoid [29].
Despite the large body of work addressing artifact-free rendering of 3D Gaussians, many of the aforementioned works imply significant performance penalties, still exhibit errors in extreme configurations or focus only on a single artifact. To the best of our knowledge, our method is the first to tackle all artifacts in a single, unified framework.
@@ 3 Method
This section covers necessary preliminary knowledge, and the components of our artifact-free 3D Gaussian renderer.
@@ 3.1 Preliminaries
Building upon the work of Kerbl et al. [13], we model the scene as a collection of 3D points, where each point is represented by an anisotropic Gaussian that serves as an approximation of the scene’s geometric structure. Each Gaussian is parameterized by a 3D mean $\bm{\upmu}\in\mathbb{R}^{3}$, a scaling factor $\mathbf{s}\in\mathbb{R}_{+}^{3}$, and a rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$. The probability density function of a Gaussian at a given position $\mathbf{x}$ is defined as
@@
This approximation introduces a projection error that leads to rendering artifacts, especially at the border of the image [12]. To circumvent these issues, Hahlbohm et al. [10] evaluate Gaussians directly in 3D. They represent the ray through each pixel $(x,y)$ as the intersection of the two planes $\bm{\uppi}_{x}=(1,0,0,-x)^{\top}$ and $\bm{\uppi}_{y}=(0,1,0,-y)^{\top}$ in screen space. They evaluate each Gaussian’s contribution along a ray, by transforming these planes into its normalized space, where the distance to the origin is equal to $\rho(\mathbf{x})$. This transformation is expressed as
@@
By computing the distance of the intersection line formed by the two planes to the origin, the Gaussian’s contribution to the pixel ray is determined. In this formulation, the integral is no longer evaluated to determine the Gaussian contribution; instead, only the maximum contribution along the ray is considered. While this approach deviates from the original 3DGS formulation [13], the training process naturally adapts to this modification. This formulation avoids the use of the inverse covariance $\mathbf{\Sigma}^{-1}$, which can become numerically instable for degenerate Gaussians (any $\mathbf{s}_{i}\approx 0$).
@@
Rendering artifacts can also arise from aliasing when the sampling rate deviates from the one used during training, particularly when the focal length or viewing distance changes (cf. Fig. 2). To address this, Yu et al. [35] propose a hybrid approach that combines a 3D smoothing filter with a 2D screen space Mip filter. The 3D smoothing filter applies a low-pass Gaussian filter to each Gaussian, where the filter size is determined by the maximum sampling frequency observed during training. The maximum sampling frequency is given by
$\hat{v}_{\text{train}}=\max\left(\left\{\hat{v}_{n}\right\}_{n=1}^{N_{c}}\right)\quad\text{, with}\quad\hat{v}=\frac{f}{d}\,\text{,}$ (6)
where $N_{c}$ is the number of training cameras, $f$ denotes the focal length in pixels and $d$ represents the $z$-component of the Gaussian’s mean $\bm{\upmu}$ in view space. The smoothing filter is defined by covariance $\hat{\mathbf{\Sigma}}=\mathbf{\Sigma}+\nicefrac{{k}}{{\hat{v}^{2}}}\ \mathbf{I}$, where $k$ controls the size of the filter. The smoothed Gaussian is then given by
@@
The 3D smoothing filter mitigates high-frequency sampling artifacts and is inherently incorporated into the Gaussian representation after training. To accommodate lower sampling rates—such as those resulting from an increased camera distance—the screen space dilation filter employed by Kerbl et al. [13] is replaced with a 2D Mip filter. This filter effectively approximates a 2D box filter in image space, replicating the physical imaging process.
@@
To optimize rendering efficiency, 3D Gaussian Splatting partitions the image into tiles. Each Gaussian is evaluated for overlap with these tiles and assigned accordingly, ensuring that each tile processes only the Gaussians that intersect with it during rendering. To determine overlap, Kerbl et al. [13] compute a screen space bounding box based on the eigenvalues of the 2D covariance matrix.
@@
Figure 3: Improvement due to our adaptive 3D dilation filter: Whereas previous methods adjust amplitudes based solely on the change of volume, often leading to excessive transparency (top left), our method adapts based on the area perpendicular to the viewing ray.
@@
To address this limitation, we replace the 2D screen space Mip filter by a full 3D filtering approach that seamlessly integrates with 3D evaluation methods, effectively preventing low-frequency aliasing. A naïve approach would be to recompute the 3D smoothing filter for each rendering view, but this proves insufficient, as it causes Gaussians to become overly transparent (cf. Fig. 3). This issue arises because the amplitude in Eq. 7 decreases according to the change in volume, while Gaussians are evaluated only at their point of maximum contribution along a ray $\mathbf{d}$, rather than being fully integrated along the ray. Consequently, incorporating the scaling change along the ray overestimates amplitude scaling for highly anisotropic Gaussians, which leads to an excessively small normalization factor
@@
in Eq. 7. Notably, this reduction occurs even when the scale change perpendicular to $\mathbf{d}$ is minimal, highlighting the need for a more robust filtering approach. Therefore, we reformulate the normalization to only factor in the change of scale perpendicular to $\mathbf{d}$ (cf. Fig. 3), that is
@@
where $\mathbf{d}$ is the normalized vector between $\bm{\upmu}$ and the camera origin $\mathbf{o}$, and $\mathbf{\Sigma}_{\perp}$ denotes the $2\times 2$ covariance matrix projected onto the subspace orthogonal to $\mathbf{d}$. It can be shown (cf. the supplementary material for a derivation) that the perpendicular scaling factor is given by
@@
where $\mathbf{d}^{\prime}=\mathbf{R}^{\top}\mathbf{d}$. Using the determinant factorization from Eq. 8, we can simplify Eq. 10 to
@@
where the updated scaling factors are given by $\mathbf{\hat{s}}_{i}=\mathbf{s}^{2}_{i}+\nicefrac{{k}}{{\hat{v}^{2}}}$. This formulation enables efficient computation while avoiding explicit matrix inversion, ensuring numerical stability. To address artifacts when moving the camera close to a Gaussian, we integrate the 3D smoothing filter of Yu et al. [35] with our proposed 3D kernel: We store the maximum sampling frequency observed across all training cameras $\hat{v}_{\text{train}}$. During rendering, the effective sampling frequency is then defined as
@@ 3.3 Perspective Correct Bounding
and a bounding step to the range $[-(\frac{\pi}{2}-\epsilon),\frac{\pi}{2}-\epsilon]$ (with a small ϵ+italic-$\epsilon\in\mathbb{R}_{+}$) to translate those bounds to the screen
@@ 4.1 Image Metrics
poppingx aliasingx distortion
@@
We begin by evaluating standard datasets, which represent in-distribution camera and view parameters. As shown in Tab. 1, our method outperforms others in nearly all metrics and matches MCMC in overall quality—while suffering from none of the artifacts. Notably, our approach prevents the optimizer from "cheating per-view" inconsistencies via popping which explains the slightly lower PSNR in Tanks & Temples. In this dataset, MCMC relies on distorted Gaussians and popping to compensate for large exposure changes, effectively "faking" full-screen adjustments. However, as shown in Fig. 6, this strategy breaks down for out-of-distribution viewing configurations.
@@
We analyze the impact of our components in Tab. 2. While standard image metrics remain largely unaffected for in-distribution test views, disabling individual components leads to distinct artifacts. Our anti-aliasing ensures stability across resolution changes and distance to the observed scene content (Tab. 4). Removing hierarchical per-pixel sorting introduces popping artifacts and allows the method to cheat with a view-inconsistent representation (again see Radl et al. [27]). Finally, disabling 3D evaluation results in projection artifacts, particularly noticeable when increasing the field of view (Fig. 6).
Dataset Mip-NeRF 360 Tanks & Temples Deep Blending
@@
Table 2: Ablation study on the effect of our individual contributions for in-distribution views. Removing individual features may actually increase image metrics, as methods can better overfit to the data set views. See the out-of-distribution evaluation for the benefits of our contributions.

### v2

@@ 1 Introduction
3D Gaussian Splatting (3DGS) [14] has recently revolutionized inverse rendering by enabling fast, differentiable rasterization of 3D Gaussian point clouds. While 2D splat evaluation is highly efficient, the projection from 3D to 2D Gaussians remains an approximation, introducing artifacts that become particularly noticeable under non-standard camera settings, such as a wide field-of-view in virtual reality rendering. Additionally, 3DGS further approximates the rendering of 2D splats by assuming them to be parallel to the current view-plane, leading to blend order inconsistencies and popping artifacts under simple camera rotations [28].
@@
Several works attempt to bridge the gap between 2D splatting and 3D ray tracing by first bounding 3D Gaussians in screen space and then computing their contributions per ray in 3D [28, 37, 11, 34, 32]. While these hybrid 2D/3D approaches address certain limitations of 3DGS, their reliance on screen space computations still makes them prone to artifacts, particularly when rendering out-of-distribution camera poses—i.e., viewpoints or parameters significantly different from the training data (cf. Fig.˜1): (1) Even when considering the highest contribution point in 3D [37, 28] or adjusting transformations per Gaussian [34], distortions may still occur. (2) Zooming in or out beyond the typical training views introduces artifacts, especially when evaluating Gaussians in 3D, as 2D anti-aliasing techniques can no longer be applied [36]. (3) Despite the use of 3D bounding planes, screen space computations become unstable when Gaussians extend behind the image plane, leading to artifacts at image boundaries [11]. (4) Simplified per-pixel sorting strategies that focus on high-opacity Gaussians, can result in inaccurate renderings, particularly for viewpoints far from the training distribution [11].
@@
Overall, our method achieves state-of-the-art reconstruction quality, while enabling artifact-free rendering in real-time. Our training and rendering source code is publicly available at https://github.com/DerThomy/AAA-Gaussians.
@@ 2 Related Work
In this section, we cover recent radiance field and 3D Gaussian Splatting methods, with a focus on artifact-free rendering of 3D Gaussian representations.
@@ 2.1 Radiance Fields & 3D Gaussian Splatting
Radiance fields have attracted widespread interest in novel view synthesis since the publication of Neural Radiance Fields (NeRF) [22], which use large coordinate-based MLPs to model view-dependent color and density. Follow-up work includes improvements in terms of anti-aliasing [1, 3], extending NeRFs to unbounded scenes [2, 26], more efficient encodings [25], or fast rendering [10, 7, 31]. Despite the aforementioned improvements, most NeRF methods still require multiple costly MLP evaluations for each pixel, manifesting in long training and rendering times.
More recently, 3D Gaussian Splatting (3DGS) [14] exploded in popularity, replacing the slower implicit representations of NeRFs with an explicit 3D Gaussian point cloud representation that can be efficiently rasterized through elliptical weighted average (EWA) splatting [39]. The initial sparse point cloud can either be initialized randomly, from Structure-from-Motion [29], or from other guidances (e.g., a pretrained NeRF [27]). To achieve good scene coverage, this point cloud has to be densified by adding, cloning, or pruning points based on screen space gradients [14], per-view saliency maps [20], depth supervision [15], or by relocating low-opacity Gaussians [16]. Additionally, strategies such as opacity decay [28] and iterative pruning [9, 8] are employed to reduce the size of the resulting point cloud.
@@ 2.2 Artifacts in 3D Gaussian Splatting
The original 3DGS suffers from a number of artifacts (underlined), which have been addressed in recent related work. Aliasing artifacts during rasterization can be solved by approximating the integral over each pixel [18], using multi-scale 3D Gaussians [35], or by applying a smoothing filter to both the 3D Gaussian and its projected 2D splat [36]. Popping artifacts occur due to the global sort of primitives before rasterization, leading to sudden changes in the blending order.
StopThePop [28] resolves this by computing Gaussian depth values along each view ray and using hierarchical $k$-buffers to improve the sort order. Other recent work relies on hybrid transparency [21], where only important contributors are sorted, while low-opacity Gaussians are blended in a "tail" [11]. Even perfect per-pixel sorting exhibits blending artifacts, due to the non-overlapping assumption of primitives, however, this cannot be solved analytically for Gaussians [19] and requires expensive volumetric integration [6, 5]. Furthermore, projection artifacts occur due to the affine approximation when projecting 3D Gaussians to 2D splats, resulting in disturbing cloud-like artifacts and elongated Gaussians, especially at the image border and for large field of view settings (e.g. in virtual reality [33]). Recent methods solve this by projecting Gaussians onto the unit sphere tangent plane [13] or using Unscented Transform [34]. Another way to avoid projection artifacts is to evaluate Gaussians in 3D. Several methods use ray tracing for this purpose [24, 19, 4], however, these methods are computationally expensive and require additional acceleration data structures, as well as specialized ray tracing hardware to achieve competitive performance. To circumvent this, several methods propose to bound the Gaussian on screen, and then evaluate it in 3D by finding its point of maximum contribution along the ray. However, these methods rely on inaccurate 2D bounds from the approximate affine projection [37, 28, 32] or Unscented Transform [34]. Only recently, Hahlbohm et al. [11] computed accurate 2D bounds through plane fitting to the 3D Gaussian ellipsoid [30].
Despite the large body of work addressing artifact-free 3D Gaussians rendering, many of the aforementioned works imply significant performance penalties, still exhibit errors in extreme configurations or focus only on a single artifact. To the best of our knowledge, our method is the first to tackle all artifacts in a single, unified framework.
@@ 3 Method
This section covers necessary preliminary knowledge, and the components of our artifact-free 3D Gaussian renderer. Our rasterizer utilizes hierarchical sorting [28] and 3D Gaussian evaluation through screen space planes [11] to remove popping and projection artifacts. Sec. 3.2 introduces our novel adaptive 3D filter that removes aliasing artifacts; Sec. 3.3 contains our perspective correct bounding approach that prevents pop-in of Gaussians from outside the view-frustum; Finally, Sec. 3.4 proposes our novel frustum-based culling, a crucial component that improves both performance and sort-order through hierarchical culling.
@@ 3.1 Preliminaries
Following Kerbl et al. [14], we model the scene as a collection of anisotropic 3D Gaussians that serves as an approximation of the scene’s geometric structure. Each Gaussian is parameterized by its 3D mean $\bm{\upmu}\in\mathbb{R}^{3}$, scaling factor $\mathbf{s}\in\mathbb{R}_{+}^{3}$, and rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$. The probability density function of a Gaussian at a position $\mathbf{x}\in\mathbb{R}^{3}$ is defined as
@@
This approximation introduces a projection error that leads to rendering artifacts, especially at image borders [13]. To circumvent these issues, Hahlbohm et al. [11] evaluate Gaussians directly in 3D: They represent the ray through each pixel $(x,y)$ as the intersection of the two screen space planes $\bm{\uppi}_{x}=(1,0,0,-x)^{\top}$ and $\bm{\uppi}_{y}=(0,1,0,-y)^{\top}$, which are then transformed into each Gaussian’s normalized space, where the distance to the origin is equal to $\rho(\mathbf{x})$:
@@
By computing the distance of the intersection line formed by the two planes to the origin, the Gaussian’s contribution to the pixel ray is determined. In this formulation, the integral is no longer evaluated to determine the Gaussian contribution; instead, only the maximum contribution along the ray is considered.
This formulation avoids the use of the inverse covariance $\mathbf{\Sigma}^{-1}$, which can become numerically instable for degenerate Gaussians (any $\mathbf{s}_{i}\approx 0$).
@@
Aliasing artifacts arise when the sampling rate during rendering deviates from the one used during training, e.g. changes in image resolution, focal length, or viewing distance (cf. Fig.˜2). To address this, Yu et al. [36] combine a 3D smoothing filter with a 2D screen space Mip filter. The 3D smoothing filter prevents high-frequency sampling artifacts by applying a low-pass Gaussian filter to each Gaussian, where the filter size is determined by the maximum sampling frequency observed during training across all $N_{c}$ training cameras, given by
$\hat{v}_{\text{train}}=\max\left(\left\{\hat{v}_{n}\right\}_{n=1}^{N_{c}}\right),\quad\text{with}\quad\hat{v}=\frac{f}{d}\,\text{,}$ (6)
where $f$ denotes the camera’s focal length in pixels and $d$ represents the $z$-component of the Gaussian’s mean $\bm{\upmu}$ in view space. The smoothing filter is defined by covariance $\hat{\mathbf{\Sigma}}=\mathbf{\Sigma}+\nicefrac{{k}}{{\hat{v}^{2}}}\ \mathbf{I}$, where $k$ controls the size of the filter. The smoothed Gaussian is then given by
@@
To prevent low-frequency sampling artifacts—e.g. when increasing the camera distance—they replace the screen space dilation filter of Kerbl et al. [14] with a 2D Mip filter, which approximates a 2D box filter in image space.
@@
To optimize rendering efficiency, 3DGS partitions the image into tiles. Each Gaussian is evaluated for overlap with these tiles and assigned accordingly, ensuring that each tile processes only the Gaussians that intersect with it during rendering. To determine overlap, Kerbl et al. [14] compute a screen space bounding box based on the eigenvalues of the 2D covariance matrix.
@@
Figure 3: Improvement due to our adaptive 3D dilation filter: Whereas previous methods adjust amplitudes based solely on the change of volume of the 3D Gaussian, often leading to excessive transparency (top left), our method adapts based on the area perpendicular to the viewing ray.
@@
To address this limitation, we replace the 2D screen space Mip filter by a full 3D filtering approach that seamlessly integrates with 3D evaluation methods, effectively preventing low-frequency aliasing. A naïve approach would be to recompute the 3D smoothing filter for each rendering view, but this proves insufficient, as it causes Gaussians to become overly transparent (cf. Fig.˜3). This issue arises because the amplitude in Eq.˜7 decreases according to the change in volume, while Gaussians are evaluated only at their point of maximum contribution along a ray $\mathbf{d}$, rather than being fully integrated along the ray. Consequently, incorporating the scaling change along the ray overestimates amplitude scaling for highly anisotropic Gaussians, which leads to an excessively small normalization factor
@@
in Eq.˜7. Notably, this reduction occurs even when the scale change perpendicular to $\mathbf{d}$ is minimal, highlighting the need for a more robust filtering approach. Therefore, we reformulate the normalization to only factor in the change of scale perpendicular to $\mathbf{d}$ (cf. Fig.˜3), that is
@@
where $\mathbf{d}$ is the normalized vector between $\bm{\upmu}$ and the camera origin $\mathbf{o}$, and $\mathbf{\Sigma}_{\perp}$ denotes the $\times 2$ 2 covariance matrix projected onto the subspace orthogonal to $\mathbf{d}$. It can be shown (cf. the supplementary material for a derivation) that the perpendicular scaling factor is given by
@@
where $\mathbf{d}^{\prime}=\mathbf{R}^{\top}\mathbf{d}$. Using the determinant factorization from Eq.˜8, we can simplify Eq.˜10 to
@@
where the updated scaling factors are given by $\mathbf{\hat{s}}_{i}=\mathbf{s}^{2}_{i}+\nicefrac{{k}}{{\hat{v}^{2}}}$. This formulation enables efficient computation while avoiding explicit matrix inversion, ensuring numerical stability. To address artifacts when moving the camera close to a Gaussian, we integrate the 3D smoothing filter of Yu et al. [36] with our proposed 3D kernel: We store the maximum sampling frequency observed across all training cameras $\hat{v}_{\text{train}}$. During rendering, the effective sampling frequency is defined as
@@ 3.3 Perspective Correct Bounding
and a bounding step to the range $[-(\frac{\pi}{2}-\epsilon),\frac{\pi}{2}-\epsilon]$ (with a small $\epsilon\in\mathbb{R}_{+}$) to translate those bounds to the screen
@@ 4.1 Image Metrics
poppingx
aliasingx
distortion
@@
We begin by evaluating standard datasets, which represent in-distribution camera and view parameters. As shown in Tab.˜1, our method outperforms others in nearly all metrics and matches MCMC in overall quality—while suffering from none of the artifacts. Notably, our approach prevents the optimizer from "cheating per-view" inconsistencies via popping which explains the slightly lower PSNR in Tanks & Temples. In this dataset, MCMC relies on distorted Gaussians and popping to compensate for large exposure changes, effectively "faking" full-screen adjustments. However, as shown in Fig.˜6, this strategy breaks down for out-of-distribution viewing configurations.
@@
We analyze the impact of our components in Tab.˜2. While standard image metrics remain largely unaffected for in-distribution test views, disabling individual components leads to distinct artifacts. Our anti-aliasing ensures stability across resolution changes and distance to the observed scene content (Tab.˜4). Removing hierarchical per-pixel sorting introduces popping artifacts and allows the method to cheat with a view-inconsistent representation (again see Radl et al. [28]). Finally, disabling 3D evaluation results in projection artifacts, particularly noticeable when increasing the field of view (Fig.˜6). Frustum-based culling has no impact on reconstruction quality, but it plays a significant role in improving performance (cf. Tab. 5).
poppingx
aliasingx
distortion
Mip-NeRF 360 Tanks & Temples Deep Blending
@@
Table 2: Ablation study on the effect of our components for in-distribution views. Removing individual features may actually increase image metrics, as method can better overfit to the data set views. See the out-of-distribution evaluation for the benefits of our contributions.